# ruledlabelfont=normalfont,labelsep=colon,strut=off

### v2

@@ 1 Introduction
Deep neural networks (DNNs) have demonstrated unprecedented capabilities in solving complex problems and handling large-scale data sets [1, 2, 3], with large models [4, 5, 6] based on these networks becoming the forefront of artificial intelligence research. While striving for high performance, fairness and bias in models have emerged as key factors affecting the practical application of artificial intelligence [7, 8, 9, 10]. Although previous studies have tended to attribute model bias to imbalances in sample sizes [11, 12, 13, 14, 15], recent research [16, 17, 18] has revealed that biases still exist in models even when sample sizes are balanced (see Fig.1a). This indicates that deeper factors are affecting the fairness of models. Therefore, exploring other intrinsic properties of data that affect model fairness, and developing more universal and comprehensive methods for predicting model biases, constitute an important and challenging task that is crucial for enhancing the fairness of models.
@@
Natural images usually follow the manifold distribution law: data of one category typically cluster near a low-dimensional manifold embedded within a high-dimensional space [19, 20, 21]. The data manifold formed by the image embeddings generated within the internal layers of DNNs is called the perceptual manifold [18]. If the collection of images for each category is considered a data manifold, then the process of data classification can be understood as effectively untangling, reducing dimensions, and separating these perceptual manifolds corresponding to different categories through the layers of the DNN [18, 22] (as shown in Fig.1b and 1c). From this geometric perspective, we hypothesize that if the data manifolds corresponding to certain categories are more complex geometrically, the DNN may encounter greater challenges in simplifying and reducing the dimensions of these perceptual manifolds, resulting in the perceptual manifolds produced by the last layers of the DNN still having relatively large intrinsic dimension (ID) [23, 24]. The differences in the IDs of perceptual manifolds across categories could lead to inconsistencies in the recognition capabilities of the classifiers at the end of the DNN for different categories, thereby affecting the fairness of the model.
@@
Estimation of the Intrinsic Dimension
@@
In this study, apart from using $ID(Z)=\frac{(tr(ZZ^{T}))^{2}}{tr((ZZ^{T})^{2})}$ [26] to estimate the intrinsic dimension of the perceptual manifold in intrinsic dimension regularization, TLE was employed in all other experiments to estimate the intrinsic dimensiona of the perceptual manifold. We will provide the code for implementing intrinsic dimension estimation at https://github.com/mayanbiao1234/Geometric-metrics-for-perceptual-manifolds.
Results
@@
Perceptual Manifold in Deep Neural Network
@@
The Intrinsic Dimension of Perceptual Manifold Exhibits Negative Correlation with DNN Performance
@@
Learning Reduces the Intrinsic Dimension
@@
ID Imbalance between Class Perceptual Manifolds Significantly Correlates with Model Fairness
@@
How Learning Affects Intrinsic Dimensional Imbalance
@@
Balancing the Intrinsic Dimensions of Class Perceptual Manifolds
@@
During model training, $L_{ID}$ can serve as a supplementary term to the existing loss function, typically formulated as $L=L_{original}+\frac{\log_{\alpha}epoch}{(\frac{L_{ID}^{smooth}}{L_{original}}).detach()}\times L_{ID}^{smooth}$, where $\alpha>1$. The purpose of the $(\frac{L_{ID}^{smooth}}{L_{original}}).detach()$-term is to keep the values of $L_{ID}$ and $L_{original}$ at the same scale. By adjusting $\log_{\alpha}epoch$, the trade-off between $L_{ID}$ and $L_{original}$ can be controlled, with the influence of $L_{ID}$ gradually increasing with the increase in epochs. The parameter $\alpha$ determines the rate of growth of $L_{ID}$. When $\alpha$ equals the epoch, $\log_{\alpha}epoch$ equals $1$. Therefore, a smaller $\alpha$ can cause the weight $\log_{\alpha}epoch$ to reach $1$ earlier. However, applying $L_{ID}$ in practical applications faces a challenge: the IDs of class perceptual manifolds keeps changing continuously during training, thus requiring real-time updates. Re-extracting embeddings of the entire dataset at each iteration not only interrupts training but also greatly increases time costs. Feature slow shift phenomenon [42, 17] indicates that as training progresses, the shift of sample embeddings becomes increasingly small, to the extent that historical versions of embeddings can approximate the current version. Inspired by this observation, we devise an end-to-end training scheme to apply IDR, as visualized in Fig.5b. The key steps of this scheme are summarized as follows:
@@
Fig.5d illustrates the impact of different values of $\alpha$ on model performance. It can be observed that increasing $L_{ID}$ in the mid-to-late stages of training can more effectively enhance the overall performance of the model. We validated the effectiveness of IDR on four benchmark image datasets [30, 31, 32] using DNN models based on both convolutional and transformer architectures [27, 28, 29]. Fig.5c demonstrates the standard deviation of class accuracy on the test set before and after employing IDR. It can be observed that the introduction of $L_{ID}$ significantly improves the fairness of deep neural networks. Additionally, $L_{ID}$ comprehensively enhances the performance of multiple baseline models, particularly achieving good results on datasets with balanced sample sizes such as CIFAR-10, CIFAR-100, and Mini-ImageNet (see Fig.5e). This performance is exciting because in the past, researchers have not paid much attention to model fairness on datasets with balanced sample sizes. In conclusion, this work provides a novel geometric perspective and tool for studying the fairness of deep neural networks, prompting AI researchers to pay attention to a broader range of imbalance phenomena.
Conclusion and Future Works

### v3

@@ 1 Introduction
Deep neural networks (DNNs) have demonstrated unprecedented capabilities in solving complex problems and handling large-scale data sets [1, 2, 3], with large models [4, 5, 6] based on these networks becoming the forefront of artificial intelligence research. While striving for high performance, fairness and bias in models have emerged as key factors affecting the practical application of artificial intelligence [7, 8, 9]. Although previous studies have tended to attribute model bias to imbalances in sample sizes [10, 11, 12, 13], recent research [14, 15, 16] has revealed that biases still exist in models even when sample sizes are balanced (see Fig.1a). This indicates that deeper factors are affecting the fairness of models. Therefore, exploring other intrinsic properties of data that affect model fairness, and developing more universal and comprehensive methods for predicting model biases, constitute an important and challenging task that is crucial for enhancing the fairness of models.
@@
Natural images usually follow the manifold distribution law: data of one category typically cluster near a low-dimensional manifold embedded within a high-dimensional space [17, 18, 19]. The data manifold formed by the image embeddings generated within the internal layers of DNNs is called the perceptual manifold [16]. If the collection of images for each category is considered a data manifold, then the process of data classification can be understood as effectively untangling, reducing dimensions, and separating these perceptual manifolds corresponding to different categories through the layers of the DNN [16, 20] (as shown in Fig.1b and 1c). From this geometric perspective, we hypothesize that if the data manifolds corresponding to certain categories are more complex geometrically, the DNN may encounter greater challenges in simplifying and reducing the dimensions of these perceptual manifolds, resulting in the perceptual manifolds produced by the last layers of the DNN still having relatively large intrinsic dimension (ID) [21, 22]. The differences in the intrinsic dimensions of perceptual manifolds across categories could lead to inconsistencies in the recognition capabilities of the classifiers at the end of the DNN for different categories, thereby affecting the fairness of the model.
@@
2 Estimation of the Intrinsic Dimension
@@
In this study, we employed two intrinsic dimension (ID) estimation methods: TLE and $ID(Z)=\frac{(tr(ZZ^{T}))^{2}}{tr((ZZ^{T})^{2})}$ [24]. While TLE was used to estimate the intrinsic dimension of perceptual manifolds in most experiments, $ID(Z)$ was specifically adopted within the intrinsic dimension regularization. Although TLE and $ID(Z)$ differ in their computational approach, both methods fundamentally aim to quantify the geometric complexity of data distributions in high-dimensional space. TLE is particularly effective in providing a precise estimation of local geometric structures; however, it involves non-linear and non-smooth operations, leading to relatively low computational efficiency and insufficient differentiability. These characteristics make it unsuitable for direct inclusion in the loss function during the optimization process. Conversely, $ID(Z)$ offers a simpler and smoother design, making it more appropriate for use in intrinsic dimension regularization during optimization. The decision to employ these two distinct estimation methods allows us to balance accuracy in experimental analysis with differentiability requirements in model optimization. We will provide the code for implementing intrinsic dimension estimation at https://github.com/mayanbiao1234/Geometric-metrics-for-perceptual-manifolds.
3 Results
@@
3.1 Perceptual Manifold in Deep Neural Network
@@
3.2 The Intrinsic Dimension of Perceptual Manifold Exhibits Negative Correlation with DNN Performance
@@
3.3 Learning Reduces the Intrinsic Dimension
@@
3.4 ID Imbalance between Class Perceptual Manifolds Significantly Correlates with Model Fairness
@@
3.5 How Learning Affects ID Imbalance
@@
3.6 Balancing the ID of Class Perceptual Manifolds
@@
During model training, $L_{ID}$ can serve as a supplementary term to the existing loss function, typically formulated as $L=L_{original}+\frac{\log_{\alpha}epoch}{(\frac{L_{ID}^{smooth}}{L_{original}}).detach()}\times L_{ID}^{smooth}$, where $\alpha>1$. The purpose of the $(\frac{L_{ID}^{smooth}}{L_{original}}).detach()$-term is to keep the values of $L_{ID}$ and $L_{original}$ at the same scale. By adjusting $\log_{\alpha}epoch$, the trade-off between $L_{ID}$ and $L_{original}$ can be controlled, with the influence of $L_{ID}$ gradually increasing with the increase in epochs. The parameter $\alpha$ determines the rate of growth of $L_{ID}$. When $\alpha$ equals the epoch, $\log_{\alpha}epoch$ equals $1$. Therefore, a smaller $\alpha$ can cause the weight $\log_{\alpha}epoch$ to reach $1$ earlier. The regularization term form proposed in [13] has proven effective, so we adopt a consistent formulation. However, the primary distinction lies in the optimization objectives: the objective in [13] is to reduce and balance the curvature of perceptual manifolds, whereas the objective of this work is to reduce and balance their intrinsic dimensions. Additionally, to address the substantial memory requirements imposed by the training framework in [13], we propose approximating the global covariance matrix with multiple local covariance matrices, thereby eliminating the need to store image embeddings and significantly reducing storage costs.
As training progresses, the IDs of class perceptual manifolds continually shift, necessitating real-time updates. Re-extracting embeddings of the entire dataset at each iteration not only interrupts training but also greatly increases time costs. Feature slow shift phenomenon [41, 15] indicates that as training progresses, the shift of sample embeddings becomes increasingly small, to the extent that historical versions of embeddings can approximate the current version. Inspired by this observation, we devise an end-to-end training scheme to apply IDR, as visualized in Fig.5b. The key steps of this scheme are summarized as follows:
@@
Fig.5d illustrates the impact of different values of $\alpha$ on model performance. It can be observed that increasing $L_{ID}$ in the mid-to-late stages of training can more effectively enhance the overall performance of the model. We validated the effectiveness of IDR on four benchmark image datasets [28, 29, 30] using DNN models based on both convolutional and transformer architectures [25, 26, 27]. Fig.5c demonstrates the standard deviation of class accuracy on the test set before and after employing IDR. Fig.6 further illustrates the per-class accuracy with and without IDR, where it is evident that incorporating $L_{ID}$ significantly reduces the bias in deep neural networks. Additionally, $L_{ID}$ comprehensively enhances the performance of multiple baseline models, particularly achieving good results on datasets with balanced sample sizes such as CIFAR-10, CIFAR-100, and Mini-ImageNet (see Fig.5e).
Compared to existing methods aimed at mitigating model bias, such as Focal Loss [42] and SeeSaw Loss [43], our method consistently achieves superior overall performance across four class-balanced datasets. Notably, while Focal Loss and SeeSaw Loss only marginally reduce model bias, IDR demonstrates a clear advantage. Importantly, Focal Loss and SeeSaw Loss differ fundamentally from our approach, as they adjust sample weights based on model predictions, while our approach derives class difficulty directly from the data, providing a new perspective and optimization target for bias mitigation. Unlike Focal Loss and SeeSaw Loss, which offer no new optimization targets, IDR introduces a novel objective by focusing on class-level intrinsic dimensions. This performance of IDR is exciting because in the past, researchers have not paid much attention to model fairness on datasets with balanced sample sizes. In conclusion, this work provides a novel geometric perspective and tool for studying the fairness of deep neural networks, prompting AI researchers to pay attention to a broader range of imbalance phenomena.
Figure 6: Per-class accuracy across four datasets using ResNet-18 as the backbone network and cross-entropy as the classification loss, comparing results before and after applying IDR.
4 Conclusion and Future Works